OpenStudy (anonymous):

The area of a square is 36m square . What is the perimeter of the square ?

OpenStudy (anonymous):

Can you help me out

OpenStudy (anonymous):

I'm guessing it's 9

OpenStudy (anonymous):

Cause 9×4=36

OpenStudy (usukidoll):

ugh I think I'm misunderstanding this formula one sec oh damn it a is the side .... sorry sorry ^_^

OpenStudy (usukidoll):

P = 4a or P = a+a+a+a

OpenStudy (anonymous):

Kk

OpenStudy (usukidoll):

AH! try letting P = 36m 36m=4a and solve for a I guess it was done already 9m = a so 9m for all four sides 36m = 9m+9m+9m+9m 36m= 36m

OpenStudy (anonymous):

Okay I'll do that

OpenStudy (anonymous):

The answer choices are 24m 6m 144m 9m

OpenStudy (usukidoll):

try 9m

OpenStudy (anonymous):

Yeah its the correct one right ?

OpenStudy (usukidoll):

yeah perimeter of the square = 4 sides P = a+a+a+a 9m x 4 = 36m. it should work.

OpenStudy (anonymous):

Yeah I agree thanks :) 😁😊 I really appreciate it